OpenStudy (anonymous):

Solve this system of by rewriting the statements into a set of 2 equations. Then solve the equations using the addition method. Question: The sum of two numbers is 21. The difference of the two numbers is 19. What are the two numbers? PLEASE HELP SOMEONE!!!!!

OpenStudy (jim766):

what does sum mean?

OpenStudy (jim766):

sum means add two numbers, but dont know what they are so call one x and the other one y x + y = 21 The other equation will be the difference of 2 numbers (x and y again) = 19 can you write that one?

OpenStudy (anonymous):

I honestly don't know what's its asking im guessing it wants to know the x and they y values im so lost

OpenStudy (jim766):

It is asking you to write two equations that represent the situation, first we need to translate x + y = 21 the sum of 2 number is 21 x - y = 19 the difference of 2 number is 19 do you see where I got that?

OpenStudy (anonymous):

i think i am making this really hard on myself but thank you for helping me through(:

OpenStudy (jim766):

now we need to solve it. this is a system of equations, we need to eliminate one of the variables. if we add the two equations together --- x + y = 21 x - y = 19 __________ 2x = 40 now that is an equation we can solve, divide both sides by x = 20 2 do you follow me at all?

OpenStudy (anonymous):

OOOOOOOOOO My gosh!!!!!! Blonde moment i finally understand it i have to set up the equations after that slove them and then i get my X and Y value Right!!

OpenStudy (jim766):

one of the number is 20, do you know how to find the other?

OpenStudy (anonymous):

um i think im not that sure :/

OpenStudy (jim766):

x = 20 , substitute that back into either one of the equations ie x + y = 21 20 + y = 21 solve that to find the other number

OpenStudy (anonymous):

Ahhh gracias(: u da best(:

OpenStudy (jim766):

thanks

OpenStudy (anonymous):

No Thank You!!!